# Better Together? A Field Experiment on

### v2

@@ 7 Generalizability and External Validity
This section identifies decision environments where our findings are most likely to be informative. We first develop a theoretical model that helps rationalize core results. We then highlight two contextual features that may have been especially important given prior research. We conclude with a cost-benefit analysis.
@@ 7.1 Model of Decisions with Algorithm Support
We develop a theoretical model to clarify how an algorithm support tool could affect decision-making (Appendix D). In line with state mandates, we model workers as conducting their best possible assessment of whether the noisy signal of child risk they perceive warrants an investigation. Children are harmed when workers underestimate risk, especially when underestimation is sizable. Under simple functional form assumptions (e.g., normally distributed noise), we obtain four key results when the algorithm reduces uncertainty in the perceived signal of risk. Improved signal could occur, for example, by allowing workers to reallocate their time and effort toward higher-order tasks (e.g., thinking rather than searching for data). The first key result of the model is that algorithm support reduces child harm, consistent with our findings in Table 2. Second, harm is especially reduced for groups with greater variance in risk (Figure 1). Third, the tool reduces incongruent overestimation of risk (Figure 2). Finally, these effects occur because workers make better assessments of child risk (Table 3 and Figure 3). Details and proofs can be found in Appendix D.
Despite its simplicity, this model captures all of our key findings. In particular, the model highlights how our results may generalize to settings where workers make informed assessments based on a noisy signal, but where large mistakes in risk assessment are particularly costly.
7.2 Leading Human Expertise
@@
7.3 Individual vs. Group Decision-Making
As required by state legislation, investigation decisions were made in small teams. Team decision-making may have contributed to the success of the algorithm tool in our setting. A longstanding literature on the “wisdom of the crowds” highlights how groups can often aggregate information more effectively than individuals (Galton, 1907). In the specific context of AI support, an emerging literature has found that groups tend to make better (Patel et al., 2019) and fairer decisions than individuals (Chiang et al., 2023).
Although we are unable to directly compare individual versus group decision-making, we can use day-to-day variation in group size to provide early empirical evidence on the impact of team size on algorithm benefits. Most teams have between four and six workers, with some outliers. Team size appears relatively balanced on observables, though we caution that team size may not be as good as random in our setting. We do not find significantly greater benefits from the algorithm for larger teams (Figure A.18). However, consistent with the literature on team size and equity, we do find greater equity gains when the algorithm was used by larger teams (Figure A.19). Though we cannot detect efficiency gains of increasing team size, our observations of team discussions suggest group social reinforcement contributed to the algorithm’s continued use and benefits, and we find greater equity gains in larger teams.
7.4 Cost-Benefit Analysis: Infinite MVPF
@@
For the MVPF numerator, we note that beneficiary willingness to pay for preventing child maltreatment is unambiguously positive. For the denominator, net government cost, we assume that the algorithm reduced the number of maltreated children by about 20, following our main estimates of 20 fewer children in the extreme right tail of the harm distribution (Table 2, Column 7). A lower-bound estimate of the public costs of maltreatment is \$62,500 per child (2020 USD) in terms of special education, criminal justice, and additional child public health expenses alone (Fang et al., 2012). The resulting estimated public savings from reduced maltreatment ($20\cdot\$62,500=\$1.25M$) are much larger than the algorithm’s implementation costs – about \$280k over two years with \$15k annual maintenance thereafter – implying an infinite MVPF of the algorithmic tool. Benefits within the trial sample alone outweigh the one-time costs of designing the tool, which has since been scaled to other agencies.
@@ 8 Conclusion
We find that humans and algorithms are complementary in our setting: both decision quality and equity improve compared to either party’s decisions (observed or hypothetical) on their own. When humans had access to algorithm support, child injury-related hospitalizations fell by 29 percent and two thirds of extreme child harm instances (top percentile of child harm index) were prevented. The algorithm was designed as an aid rather than a replacement for social workers, who were acutely aware of what information was included in the algorithm versus available exclusively to them. In this context, algorithm support helped workers parse through new, complementary information about families. The algorithm helped workers reduce their tendency to investigate Black children, particularly those scored as low risk, and conversely human oversight helped reduce CPS surveillance that the algorithm would have otherwise induced for Hispanic and low-income households. Human-algorithm complementarity is thus salient not just for efficiency but also for equity. Our model and documented mechanisms rationalize these findings and suggest they could extend to other decision settings when algorithms reduce the noise in the decision-making process, such as through prompting richer discussion of critical information or reducing extraneous cognitive load. Existing literature suggests that design features – in particular which information was available to the algorithm versus humans, and that the algorithm was deployed as part of a group discussion – may have contributed to our findings.
@@
iii Number of Re-Referrals
iv Investigator-Assigned Risk Rating
Notes: These subfigures show the relationship between the algorithmic tool risk score, binned by quintile, and other relevant outcomes. Each point represents the mean of the outcome variable in the control group at a given quintile of predicted risk, with a 90% confidence interval reported. The algorithm was trained on prior data to predict a child’s removal from their home to foster care within two years of a referral. Investigator-assigned risk ratings (Panel IV) come from visits to the home and are only available for the set of screened-in children.
Figure A.4: Algorithmic Tool is Predictive of Emergency Room Visits over Time
Control Group
Notes: This figure shows the relationship between the algorithm’s predicted risk score and emergency room visits in the state relative to the timing of a child’s first CPS appearance during the trial. Month 0 is the first month that a child is referred to CPS during the trial. Plot includes the control group only. The risk score predicts a child’s likelihood of being removed to foster care within two years of the referral. The vertical axis shows the percent of children with any emergency room visit in a given month. The solid black line shows the incidence of ER visits for children with an algorithm risk score (generated at month 0) above the median score, whereas the light gray dashed line shows the incidence of ER visits for children with a risk score below the median score.
@@
Notes: This figure presents a stylized summary of teams’ decision workflow during the trial. Algorithm scores were available for the intervention group, and unavailable for the control group. Workers were encouraged, but not required, to check the algorithm and write down “No Score” or “Tool [Score]” with the specific listed score(s).
@@
Notes: This figure presents mean child harm by algorithm score quintile, with and without the tool available, with 90% confidence intervals. The sample is restricted to children randomized during the first year of the trial to allow for at least eight months of potential hospitalization following randomization. The harm index is constructed as a standardized sum of standardized outcomes including a child’s number of high-priority hospital admissions, number of admissions with a listed injury, number of avoidable ER visits, any admission with a confirmed maltreatment code, and any visit with an intentional injury (assault, self-harm) code. The index is standardized (mean 0, variance 1) on the control group and a lower value is considered better.
Figure A.7: Little Evidence of Long-Run Worker Learning
Notes: This figure examines whether there is any evidence of long-run learning by assessing whether benefits of the algorithm for the intervention group decrease over time. On the left, we present the main (full hospital sample) point estimate of providing the algorithmic tool on child harm (“All”). On the right, we then split this sample into three equally sized groups by date of referral and estimate the effects of providing the tool on the harm index in each of these sample terciles. T1 is the first third of children referred during the trial, T2 the second, and T3 the final third of children referred. 90% confidence intervals are reported throughout. The sample is restricted to children randomized during the first year of the trial, to allow for at least eight months of potential hospitalization following randomization (hospital outcomes sample).
Figure A.8: Tool Reduces Racial Disparities for Low-Risk-Score Children
@@
Notes: This figure presents point estimates and standard errors of regressing the listed outcomes on tool availability, as in equation (1), for the full sample of children. Estimates are presented with standardized outcome variables (mean 0, variance 1) for comparability of coefficient magnitudes across different outcomes. We measure margins of CPS intervention, such as high-risk-track screen-ins, using child welfare administrative data. Additional margins of intended intervention, such as voluntary resources sent to screened-out families, are inferred from team discussion notes. Some interventions are unobserved: services paid for by Medicaid, for example, are not included in administrative data. Workers can recommend opening a prevention case (row 2), can send the referral to law enforcement (row 3), and can connect the family to other resources (row 1) instead or in addition to screening in (row 4). If workers decide to screen in a family, they also choose whether the screen-in is on a low-risk or high-risk track (rows 5 and 6). Low-risk track investigations are a type of differential response that is less intrusive for the family. Formal investigations are only run for the high-risk track. As a result, whether the investigation was founded (i.e., substantiated; row 8) is only available for the high-risk screen-in track. Row 9 shows the change in any other services paid for directly by CPS. The number of minutes it took workers to make a decision (row 7) is measured using time stamp data and is only available for observations where edits to the notes were not made after normal discussion meeting times. The final two rows (rows 10-11) report estimates on whether a child was ever removed to foster care, and the number of re-referrals. We report more details on these last two outcomes in a separate table (Appendix Table A.6). Randomization procedure controls are included throughout. Standard errors are clustered at the household level and are shown at the 90 percent confidence level.
@@
Notes: This figure presents point estimates and standard errors of regressing the listed outcomes on tool availability interacted with algorithm score for the full sample of children. We report point estimates and standard errors for the coefficient on the interaction of tool availability with algorithm score. Following our pre-registration plan, we use the maximum score of all children on the referral, as this maximum score was most salient to workers and most of the decisions listed typically apply to all children in a household. Estimates are presented with standardized outcome variables (mean 0, variance 1) for comparability of coefficient magnitudes across different outcomes. We measure margins of CPS intervention, such as high-risk-track screen-ins, using child welfare administrative data. Additional margins of intended intervention, such as voluntary resources sent to screened-out families, are inferred from team discussion notes. Some interventions are unobserved: services paid for by Medicaid, for example, are not included in administrative data. Workers can recommend opening a prevention case (row 2), can send the referral to law enforcement (row 3), and can connect the family to other resources (row 1) instead or in addition to screening in (row 4). If workers decide to screen in a family, they also choose whether the screen-in is on a low-risk or high-risk track (rows 5 and 6). Low-risk track investigations are a type of differential response that is less intrusive for the family. Formal investigations are only run for the high-risk track. As a result, whether the investigation was founded (i.e., substantiated; row 8) is only available for the high-risk screen-in track. Row 9 shows the change in any other services paid for directly by CPS. The number of minutes it took workers to make a decision (row 7) is measured using time stamp data and is only available for observations where edits to the notes were not made after normal discussion meeting times. The final two rows (rows 10-11) report estimates on whether a child was ever removed to foster care, and the number of re-referrals. We report more details on these last two outcomes in a separate table (Appendix Table A.6). Randomization procedure controls are included throughout. Standard errors are clustered at the household level and are shown at the 90 percent confidence level.
@@
Figure A.14: LLM Prediction of Algorithm Support from Discussion Text Data
Notes: This figure shows the ROC curve of a fine-tuned DistilBERT LLM predicting whether workers had algorithm support from their discussion notes. The model was trained on a randomly selected 80 percent set of referrals, and the predictions displayed in the above graph are for a random 20 percent test sample of the text discussion notes.
@@
Notes: For each child subgroup as reported in the title of each figure, we plot the health disparities child harm for three different scenarios: when humans do not have access to the algorithm (light blue short-dashed line) which is our control group, when humans have access to the algorithm (solid dark blue line, with shaded area representing a 90% confidence interval; main estimates presented in Table 2), and the best-case scenario for the algorithm (orange dash-dot line). We assume that the effects of a screen-in are homogeneous across groups. The best-case scenario for the algorithm assumes that the children who were screened in by humans, but for whom the algorithm would have screened out, would have been just as well off under the algorithm decision. Effects of the algorithm screening in children who humans screened out depends on the hypothetical benefits of a screen-in on the child harm index (horizontal axis). SNAP (food stamp) eligibility is a proxy for lower socioeconomic status.
@@
Notes: This figure shows a disproportionate improvement in algorithm treatment effects for members of disadvantaged groups with larger teams. The standardized index of child harm is regressed on a triple-interaction of algorithm availability (binary), demographic group (binary), and team size (continuous). The vertical axis indicates the coefficient estimate and standard error bounds on the triple-interaction term, and the horizontal axis indicates the demographic group that is compared to all children not in that group. “Any” consists of all children in the sample who fall into at least one of the disadvantaged categories. Randomization procedure controls are included throughout, as well as unit and month-year fixed effects. Standard errors are clustered at the household level and are shown at the 95 percent confidence level.

### v3

@@ 7 Generalizability and External Validity
This section identifies decision environments where our findings are most likely to be informative. We develop a theoretical model that helps rationalize core results, and highlight two contextual features that may have been especially important given prior research. We conclude with a cost-benefit analysis.
7.1 Group Decision-Making
Group-based decision-making is common in high-stakes settings with large amounts of information and professional discretion: for example juries and parole boards, criminal justice and health care expert boards, and hiring committees. Similarly, as required by state legislation, child protection investigation decisions in the partner agency were made in small teams. Team decision-making may have contributed to the success of the algorithm tool in our setting. A longstanding literature on the “wisdom of the crowds” highlights how groups can often aggregate information more effectively than individuals (Galton, 1907). The literature on jury size and composition indeed suggests that larger groups reach better and fairer decisions (Waters, 2004; Anwar et al., 2012). In the specific context of AI support, an emerging literature has found that groups tend to make better (Patel et al., 2019) and fairer decisions than individuals (Chiang et al., 2023).
Although we are unable to directly compare individual versus group decision-making, we can use day-to-day variation in group size to provide early empirical evidence on the impact of team size on algorithm benefits. Most teams have between four and six workers, with some outliers. Team size appears relatively balanced on observables, though we caution that team size may not be as good as random in our setting. Point estimates for benefits of the algorithm appear to be potentially greater for larger teams, but we cannot statistically distinguish between these estimates (Figure A.19). However, consistent with the literature on team size and equity, we do find greater equity gains on some margins when the algorithm was used by larger teams (Figure A.20). Though we cannot detect efficiency gains of increasing team size, our observations of team discussions suggest group social reinforcement (regardless of size) contributed to the algorithm’s continued use and benefits, and we find greater equity gains in larger teams.
@@ 7.2 Model of Decisions with Algorithm Support
We develop a theoretical model to clarify how an algorithm support tool could affect decision-making (Appendix D). In line with state mandates, we model workers as conducting their best possible assessment of whether the noisy signal of child risk they perceive warrants an investigation. Children are harmed when workers underestimate risk, especially when underestimation is sizable. Under simple functional form assumptions (e.g., normally distributed noise), we obtain four key results when the algorithm reduces uncertainty in the perceived signal of risk. Improved signal could occur, for example, by allowing workers to reallocate their time and effort toward higher-order tasks, such as thinking rather than searching for data. The first key result of the model is that algorithm support reduces child harm, consistent with our findings in Table 1. Second, harm is especially reduced for groups with greater variance in risk (Figure 1). Third, the tool reduces incongruent overestimation of risk (Figure 2). Finally, these effects occur because workers make better assessments of child risk (Table 2 and Figure 3). Details and proofs are found in Appendix D.
Despite its simplicity, this model captures all of our key findings. The model highlights how our results may generalize to settings where workers make informed assessments based on a noisy signal, but where large mistakes in risk assessment are particularly costly.
7.3 Design Choices and Human Expertise
@@
7.4 Cost-Benefit Analysis: MVPF Estimate
@@
For the MVPF numerator, we note that beneficiary willingness to pay for preventing child maltreatment is unambiguously positive. For the denominator, net government cost, we assume that the algorithm reduced the number of maltreated children by about 20, following our main estimates of 20 fewer children in the extreme right tail of the harm distribution (Table 2, Column 7). A lower-bound estimate of the public costs of maltreatment is \$62,500 per child (2020 USD) in terms of special education, criminal justice, and additional child public health expenses alone (Fang et al., 2012). [^20] We cannot assess whether our findings are from a reduction in the intensive or extensive margin of maltreatment incidence. We are unaware of any study estimating costs of maltreatment for these two margins separately and so are unable to benchmark against the cost of one additional maltreatment incident. The resulting estimated public savings from reduced maltreatment ($20\cdot\mathdollar 62,500=\mathdollar 1.25M$) are much larger than the algorithm’s implementation costs – about \$280k over two years with \$15k annual maintenance thereafter – implying an infinite MVPF of the algorithmic tool. A limitation of our approach is that many of the children involved in the trial may have already been victims of maltreatment, confounding our estimates of government savings from curtailed abuse and neglect. However – to the extent that the costs of maltreatment are cumulative [^21] Prior studies have found an association between negative outcomes and number of maltreatment incidents or number of types of experienced maltreatment (e.g., Currie and Tekin 2012), consistent with a clinically hypothesized dosage-response relationship, though causality is difficult to establish. and/or some fraction of children were not prior victims – even if only a fourth of the lower-bound estimate of government costs for child maltreatment were saved, the resulting MVPF would remain infinite, suggesting some robustness of our findings to alternative cost assessment approaches. A large share of the costs of the tool were up-front fixed costs, implying a potential increase in cost effectiveness from scaling the tool to other agencies, which has occurred since the trial.
@@ 8 Conclusion
We find that humans and algorithms are complementary in our setting: both decision quality and equity improve compared to either party’s decisions (observed or hypothetical) on their own. When humans had access to algorithm support, child injury-related hospitalizations fell by 21 percent and two thirds of extreme child harm instances (top percentile of child harm index) were prevented. The algorithm was designed as an aid rather than a replacement for social workers, who were acutely aware of what information was included in the algorithm versus available exclusively to them. In this context, algorithm support helped workers parse through new, complementary information about families. The algorithm helped workers reduce their tendency to investigate Black children, particularly those scored as low risk, and conversely human oversight helped reduce CPS surveillance that the algorithm would have otherwise induced for Hispanic and low-income households. Human-algorithm complementarity is thus salient not just for efficiency but also for equity. Our model and documented mechanisms rationalize these findings and suggest they could extend to other decision settings when algorithms reduce the noise in the decision-making process, such as through prompting richer discussion of critical information or reducing extraneous cognitive load. Existing literature suggests that design features – in particular which information was available to the algorithm versus humans, and that the algorithm was deployed as part of a group discussion (Galton, 1907; Simoiu et al., 2019; Chiang et al., 2023) – may have contributed to our findings.
@@
iii Investigator-Assigned Risk Rating
Notes: These subfigures show the relationship between relevant outcomes and the algorithm risk score, binned by quintile for visual clarity. Each point represents the mean of the outcome variable in the control group at a given quintile of predicted risk, with a 90% confidence interval reported. The algorithm was trained on prior data to predict a child’s removal from their home to foster care within two years of a referral. Investigator-assigned risk ratings (Panel iii) come from visits to the home and are only available for the set of screened-in children.
@@
Notes: This figure presents a stylized summary of teams’ decision workflow during the trial. Algorithm scores were available for the intervention group, and unavailable for the control group. Workers were encouraged, but not required, to check the algorithm and write down “No Score” if the score was unavailable or “Tool [Score]” with the specific listed score(s).
@@
Notes: This figure presents mean child harm by algorithm score quintile, with and without the tool available, with 90% confidence intervals. The sample is restricted to children randomized during the first year of the trial to allow for several months of potential hospitalization following randomization. The harm index is constructed as a standardized sum of standardized outcomes including a child’s number of high-priority hospital admissions, number of admissions with a listed injury, number of avoidable ER visits, substance exposure incidents, number of admissions with a confirmed maltreatment code, and number of visits with an intentional injury (assault, self-harm) code. The index is standardized (mean 0, variance 1) on the control group and a lower value is considered better.
Figure A.6: Bayesian Posterior Distribution of Effects on Harm Index Components
Notes: These figures report the posterior distribution of the effect of algorithm support on six components of the harm index: counts of high priority visits, injury visits, avoidable ER visits, substance exposure incidents, visits with explicit maltreatment codes (rare), and visits with intentional injuries. We assume as an uninformative uniform prior that effects could have been anywhere in the range of -100 percent to +100 percent (gray line centered around vertical line at 0). Following Brännlund et al. (2024), we then calculate the posterior distribution given the data we observe (posterior distribution outlined in blue; observed mean marked by dash-dot vertical blue line). 90% credible intervals are reported as the shaded areas.
Figure A.7: Estimated Effects by Time of First Appearance in the Trial
Notes: This figure presents estimates of the tool on harm by the time a child was first seen during the trial. On the left, we present the main (full sample) point estimate of providing the algorithm tool on child harm (“All”). The full sample is then split into four equally sized groups by date of referral and estimate the effects of providing the tool on the harm index in each of these sample quarters. Q1 is the first quarter of children referred during the trial, Q2 the second, and so on. 90% confidence intervals are reported throughout. Absence of impacts for children referred in the final quarter (Figure A.7) could be due to a limited window of time to observe hospital outcomes for this subsample during the trial. After the end of the trial, the algorithm tool became available for all children being referred, attenuating effects.
Figure A.8: Bayesian Posterior Distribution of Effects on Disparities with an Uninformative Prior
Notes: These figures report the posterior distribution of the effect of algorithm support on four harm index disparity margins: race, ethnicity, gender and socioeconomic status. We assume as an uninformative uniform prior that effects could have been anywhere in the range of -200% to +200% (gray line centered around vertical line at 0). Given the data and following Brännlund et al. (2024), we then calculate the posterior distribution given the data we observe (posterior distribution outlined in blue; observed mean marked by dash-dot vertical blue line). 90% credible intervals are reported as the shaded areas.
Figure A.9: Tool Reduced Racial Disparities for Low-Risk-Score Children
@@
Notes: This figure presents point estimates and confidence intervals of regressing the listed outcomes on tool availability, as in equation (1), for the full sample of children (unless otherwise specified). Estimates are presented with standardized outcome variables (mean 0, variance 1) for comparability of coefficient magnitudes across different outcomes. We measure margins of CPS intervention, such as high-risk-track screen-ins, using child welfare administrative data. Additional margins of intended intervention, such as voluntary resources sent to screened-out families, are inferred from team discussion notes. Some interventions are unobserved: services paid for by Medicaid, for example, are not included in administrative data. Workers can recommend opening a prevention case (row 2), can send the referral to law enforcement (row 3), and can connect the family to other resources (row 1) instead or in addition to screening in (row 4). If workers decide to screen in a family, they also choose whether the screen-in is on a low-risk or high-risk track (rows 5 and 6). Low-risk track investigations are a type of differential response that is less intrusive for the family. Formal investigations are only run for the high-risk track. As a result, having a founded (i.e., substantiated; row 8) investigation requires having a high-risk track investigation. Row 9 shows the change in any other services paid for directly by CPS. The number of minutes it took workers to make a decision (row 7) is measured using time stamp data and is only available for observations where edits to the notes were not made after normal discussion meeting times. The final two rows (rows 10-11) report estimates on whether a child was ever removed to foster care, and the number of re-referrals. We report more details on these last two outcomes in a separate table (Appendix Table A.7). Randomization procedure controls are included throughout. Standard errors are clustered at the household level and are shown at the 90 percent confidence level.
@@
Notes: This figure presents point estimates and confidence intervals of regressing the listed outcomes on tool availability interacted with algorithm score for the full sample of children (unless otherwise specified). We report point estimates and standard errors for the coefficient on the interaction of tool availability with algorithm score. Following our pre-registration plan, we use the maximum score of all children on the referral, as this maximum score was most salient to workers and most of the decisions listed typically apply to all children in a household. Estimates are presented with standardized outcome variables (mean 0, variance 1) for comparability of coefficient magnitudes across different outcomes. We measure margins of CPS intervention, such as high-risk-track screen-ins, using child welfare administrative data. Additional margins of intended intervention, such as voluntary resources sent to screened-out families, are inferred from team discussion notes. Some interventions are unobserved: services paid for by Medicaid, for example, are not included in administrative data. Workers can recommend opening a prevention case (row 2), can send the referral to law enforcement (row 3), and can connect the family to other resources (row 1) instead or in addition to screening in (row 4). If workers decide to screen in a family, they also choose whether the screen-in is on a low-risk or high-risk track (rows 5 and 6). Low-risk track investigations are a type of differential response that is less intrusive for the family. Formal investigations are only run for the high-risk track. As a result, having a founded (i.e., substantiated; row 8) investigation requires having a high-risk track investigation. Row 9 shows the change in any other services paid for directly by CPS. The number of minutes it took workers to make a decision (row 7) is measured using time stamp data and is only available for observations where edits to the notes were not made after normal discussion meeting times. The final two rows (rows 10-11) report estimates on whether a child was ever removed to foster care, and the number of re-referrals. We report more details on these last two outcomes in a separate table (Appendix Table A.7). Randomization procedure controls are included throughout. Standard errors are clustered at the household level and are shown at the 90 percent confidence level.
@@
Figure A.15: Prediction of Algorithm Support from Discussion Text Data
Using 100+ Topics Generated from Word Embeddings by BERTopic
Notes: This figure shows the three-fold cross-validation ROC curves on the trial sample at the referral level since discussion notes are the same for all children on the same referral ($N=2,668$). A small number of referrals from our sample that have no discussion notes are omitted from the analysis. The topics selected by the BERTopic model are used to predict whether workers had access to the algorithm tool for two sets of folds of the data, and are then validated on the remaining fold. The cross-validated mean AUC and cross-validated standard deviation are reported at the bottom of the figure. The BERTopic model was run on a set of eleven thousand referrals. It converts the text data using sentence transformers into word embeddings, which it then reduces dimensionality of using a state-of-the-art UMAP to avoid the curse of dimensionality when applying the HDBSCAN clustering algorithm for these embeddings. In practice, we used a seeded topic model to guide the model through some of the technical, profession-specific terminology used in the notes and ensure greater replicability across runs.
@@
Notes: For each child subgroup as reported in the title of each figure, we plot the health disparities in child harm for three different scenarios: when humans do not have access to the algorithm (light blue short-dashed line) which is our control group, when humans have access to the algorithm (solid dark blue line, with shaded area representing a 90% confidence interval; main estimates presented in Table 1), and the best-case scenario for the algorithm (orange dash-dot line). We assume that the effects of a screen-in are homogeneous across groups. The best-case scenario for the algorithm assumes that the children who were screened in by humans, but for whom the algorithm would have screened out, would have been just as well off under the algorithm decision. Effects of the algorithm screening in children who humans screened out depends on the hypothetical benefits of a screen-in on the child harm index (horizontal axis). SNAP (food stamp) eligibility is a proxy for lower socioeconomic status.
@@
Notes: This figure shows a disproportionate improvement in algorithm treatment effects for members of disadvantaged groups with larger teams. The standardized index of child harm is regressed on a triple-interaction of algorithm availability (binary), demographic group (binary), and team size (continuous). The vertical axis indicates the coefficient estimate and standard error bounds on the triple-interaction term, and the horizontal axis indicates the demographic group that is compared to all children not in that group. “Any” consists of all children in the sample who fall into at least one of the disadvantaged categories. Randomization procedure controls are included throughout, as well as unit and month-year fixed effects. Standard errors are clustered at the household level and are shown at the 90 percent confidence level.